OpenStudy (anonymous):

Lucille knocks her textbook off her desk. After it has fallen for 2 seconds, what is its velocity?

OpenStudy (shane_b):

Use the kinematic equation: \[\Delta V =at\]You know that a = 9.8 m/s^2 and t = 2 seconds.

OpenStudy (zepp):

The acceleration constant, \(a\), describes the velocity of an object after a certain time. \(9.8m/s^2\) mean that the object's velocity is about 9.8m/s, due to the gravitational force, \(\LARGE 9.8m/s^2=\frac{9.8m/s}{s}\), 9.8 meters per second PER second. Therefore, if an object is free falling and neglicting air resistantance, after 1 second, the velcity would be 9.8m/s, and 2 second, the velocity would be 9.8m/s TIMES 2 = 19.6m/s, after 3 seconds, the velocity would be 9.8m/s TIMES 3!, 29.4m/s, and so on.